# EPO: Entropy-regularized Policy Optimization for LLM Agents Reinforcement Learning

### v1

@@ 5.4 Model Study
(a) EPO-Base vs EPO-Decay
(b) EPO vs EA Success Rate
(c) Early vs Late Entropy
Figure 4: Model studies on ScienceWorld. (a) Consistent entropy regularization outperforms decaying schedules, preventing cascade failure. (b) EPO achieves near-perfect success while EA plateaus at 0.5–0.6 due to reasoning degradation. (c) Decay schedules prematurely suppress crucial early-turn exploration, triggering late-stage uncertainty propagation.
5.4.1 Study of Entropy Regularization
We compare our standard method, PPO+EPO-Base, which applies consistent entropy regularization throughout training, against PPO+EPO-Decay, a variant employing a dynamic schedule that assigns higher entropy weight during initial training phases to promote exploration and systematically reduces it in later phases to encourage exploitation. Counter-intuitively, empirical results in Figure 4 show this decay strategy consistently underperforms across all metrics. While the decay schedule successfully reduces policy entropy in later training stages, it prematurely suppresses exploration in the crucial initial turns of each episode. Panel (c) reveals that comparing average entropy of the first 10 tokens (“Early Steps”) with the last 10 tokens (“Late Steps”) shows insufficient early exploration locks agents into suboptimal strategies from which they cannot recover, even as the policy becomes more deterministic.
$\triangleright$ Insight: Directly changing loss weights fails in LLM agent scenarios due to the exploration-exploitation cascade failure in multi-turn settings. Unlike single-turn tasks, multi-turn environments exhibit strong temporal dependencies where early steps fundamentally shape late-step outcomes. The decay schedule triggers excessive early-stage exploration that creates unstable foundations, systematically committing to suboptimal behavioral patterns. These flawed early trajectories then lead to late-stage uncertainty propagation, where accumulated uncertainty compounds and prevents coherent strategy formation. Therefore, for complex multi-turn sparse-reward tasks, maintaining robust and consistent exploration pressure throughout all trajectory steps is essential to avoid cascade failure, rather than following conventional exploration-to-exploitation scheduling.
5.4.2 Study of Entropy-Shaped Advantage
We also compare our Entropy-smoothed Policy Optimization (EPO) with the Entropy-based Advantage (EA) shaping method from (Cheng et al., 2025b). As shown in Figure 4(b), while PPO+EA improves over the baseline, PPO+EPO achieves substantially superior performance in both final performance and convergence speed, reaching near-perfect success rates ($\sim$1.0) while EA plateaus at 0.5–0.6. The primary difference lies in the gradient signal and how it affects the underlying LLM’s capabilities. EA uses a detached entropy term acting as an indirect intrinsic reward, providing no gradient signal to explicitly increase entropy. In contrast, EPO integrates entropy directly into the policy loss, enabling direct gradient signals $\nabla_{\theta}L^{H}(\theta)$ to guide the policy towards more exploratory behavior. Furthermore, EA’s hard clipping on the advantage bonus can induce training instability and its myopic nature considers only instantaneous entropy.
$\triangleright$ Insight: Directly modifying the policy loss for LLM agent RL can severely impair the model’s reasoning abilities in agent scenarios—capabilities that were not developed during pretraining. Since LLMs are not pretrained on agent-specific tasks, aggressive entropy regularization directly injected into the policy loss disrupts the model’s learned representations and reasoning pathways. Our EPO method addresses this by using temporal smoothing with historical entropy windows, which preserves the LLM’s inherent reasoning capabilities while providing exploration guidance. This decoupled regularization maintains the integrity of the value signal and the pretrained knowledge, leading to more robust and effective learning without degrading the model’s fundamental abilities.
@@ 6 Conclusion
In this work, we identified and addressed the exploration-exploitation cascade failure which is a fundamental challenge unique to training LLM agents in multi-turn environments with sparse rewards. Our proposed EPO framework addresses this through trajectory-aware entropy computation, entropy smoothing regularization, and adaptive phase-based weighting—mechanisms that together prevent the dangerous entropy oscillations that destabilize training. Empirical results demonstrate up to 152% performance improvement on ScienceWorld and 19.8% on ALFWorld, transforming previously untrainable scenarios into smoothly converging optimization problems. This work establishes that multi-turn LLM agent training requires fundamentally different entropy control than traditional RL, opening new directions for developing effective training methods for LLM Agents.
Appendix A Theoretical Analysis of Entropy-Smoothed Policy Optimization
A.1 Preliminary Setup
We build upon the analytical framework from the reference work (Shen, 2025; Sutton et al., 1999; Klein et al., 2023), leveraging their established lemmas without reproducing full proofs.
Lemma 1 (Performance Difference).
For any two policies $\pi$ and $\pi^{\prime}$, and any state $s_{0}$:
$V^{\pi}(s_{0})-V^{\pi^{\prime}}(s_{0})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{H-1}A_{t}^{\pi^{\prime}}(s_{t},a_{t})\bigg|s_{0}\right]$
Lemma 2 (Entropy Gradient).
For a softmax policy $\pi_{\theta}$:
$\nabla\mathcal{H}(\pi_{\theta})=-\mathbb{E}_{\pi_{\theta}}\left[\sum_{h=0}^{H-1}\nabla\log\pi_{\theta}(a_{h}|s_{h})\sum_{t=h}^{H-1}\log\pi_{\theta}(a_{t}|s_{t})\right]$
Lemma 3 (Entropy Bias).
For the entropy-regularized objective with optimal policy $\pi_{\lambda}^{*}=\arg\max_{\pi}V_{\lambda}^{\pi}(\mathcal{D})$:
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{\frac{1}{H}}}$
where $\mathcal{A}_{H}^{*}(s_{0})$ denote the set of optimal action sequences for a given initial state $s_{0}$ and a horizon of $H$.
Lemma 4 (Performance Bound under Gradient Norm).
@@ A.1 Preliminary Setup
If $\|\nabla V_{\lambda}^{\pi_{\theta}}(\mathcal{D})\|\leq\epsilon$:
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})\leq\frac{1}{2\lambda}\frac{|\mathcal{D}|^{2}}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\epsilon^{2}$
A.2 The EPO Performance Bound
The EPO objective $V_{\lambda,\beta}^{\pi_{\theta}}$ can be understood as a three-fold optimization target: maximizing returns (exploitation), encouraging policy diversity (exploration), and constraining exploration behavior over time (stabilization). This addresses the core challenge that standard maximum entropy RL solves "whether to explore," but not "how to explore stably."
Definition 5 (EPO Objective).
The EPO-regularized objective is:
$V_{\lambda,\beta}^{\pi_{\theta}}(\mathcal{D}):=V^{\pi_{\theta}}(\mathcal{D})+\lambda\mathcal{H}(\pi_{\theta})-\lambda\beta_{k}L^{\text{smooth}}(\theta)$
where $L^{\text{smooth}}(\theta)$ is defined as in Equation 6:
@@ A.2 The EPO Performance Bound
$L^{\text{smooth}}(\theta)=\frac{1}{B}\sum_{j=1}^{B}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{|\tau_{j,t}|}\sum_{i=1}^{|\tau_{j,t}|}\mathcal{P}_{j,t,i}$
with penalty function:
$\mathcal{P}_{j,t,i}=\begin{cases}0,&\text{if }\kappa_{l}\bar{H}^{\mathcal{W}_{k}}\leq\mathcal{H}_{j,t,i}\leq\kappa_{r}\bar{H}^{\mathcal{W}_{k}}\\ \alpha,&\text{otherwise}\end{cases}$
Proposition 6 (Improved Performance Bound with EPO).
Assume the policy follows a softmax parameterization. If the EPO policy gradient satisfies $\|\nabla V_{\lambda,\beta}^{\pi_{\theta}}(\mathcal{D})\|\leq\epsilon$, then for any query $s_{0}$, the policy suboptimality is bounded by:
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\underbrace{\frac{|\mathcal{D}|^{2}}{C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})}\frac{\epsilon^{2}}{2\lambda}}_{\text{Optimization Error}}+\underbrace{\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{\frac{1}{H}}}}_{\text{Standard Entropy Bias}}-\underbrace{\lambda\beta_{k}\Phi^{\pi_{\theta},\pi^{*}}(s_{0})}_{\text{Corrective Bias from Smoothing}}$
where:
  • $C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})$ is a problem-dependent constant (specified below)
  • $\Phi^{\pi_{\theta},\pi^{*}}(s_{0}):=\mathbb{E}_{\pi_{\theta}}[L^{\text{smooth}}(\theta)|s_{0}]-\mathbb{E}_{\pi^{*}}[L^{\text{smooth}}(\theta^{*})|s_{0}]$ represents the smoothing gap
@@
Let $\pi_{\lambda,\beta}^{*}=\arg\max_{\pi}V_{\lambda,\beta}^{\pi}(\mathcal{D})$ be the optimal policy under EPO. By optimality:

### v2

@@ 5.4 Model Study
Turn-wise vs. Uniform Entropy Weighting. We compare consistent entropy regularization (PPO+EPO-Base) against a decaying schedule (PPO+EPO-Decay), which applies higher entropy weights to early turns and lower weights to later turns. Counter-intuitively, Figure 4(a) shows that decay consistently underperforms by prematurely suppressing early-turn exploration, locking agents into suboptimal strategies. Figure 4(b) further reveals the mechanism: while both methods exhibit similar early-step entropy, the decay schedule causes late-step entropy to drop too rapidly, reducing the policy’s ability to recover from suboptimal trajectories.
Decoupled Regularization vs. Advantage Reshaping. We compare EPO against the Entropy-based Advantage (EA) method from cheng2025reasoning. As shown in Figure 4(c), EPO converges to near-perfect success rates ($\sim$1.0) while EA plateaus at 0.5–0.6. EA incorporates entropy directly into the advantage function, reshaping the policy gradient. In contrast, EPO decouples entropy as a separate regularization term, providing direct gradient signals while preserving the integrity of advantage-based credit assignment.
Cumulative vs. Sliding Entropy Window. Our cumulative entropy window $\mathcal{W}_{k}$ is an intentional design choice for long-horizon training. With $K$ = 120-150 steps, the cumulative average adapts quickly when $k$ is small (weight $\sim 1/k$) and stabilizes later to anchor entropy against cascade failure. This design synergizes with the adaptive $\beta_{k}$: while $\beta_{k}$ modulates smoothing strength dynamically, a sliding window would create conflicting adaptation mechanisms.
These studies reveal two key insights: (1) multi-turn sparse-reward tasks require sustained exploration rather than conventional exploration-to-exploitation scheduling, and (2) decoupled entropy regularization with temporal smoothing is superior to advantage reshaping for maintaining long-horizon policy stability.
6 Theoretical Analysis
We provide theoretical analysis to understand EPO’s design and characterize when and why it outperforms standard entropy regularization. Our analysis reveals that EPO is a principled approach grounded in constrained optimization theory, with particular advantages in multi-turn settings where entropy instability compounds across turns. Full proofs are provided in Appendix A.
Proposition 6.1 (Exact Penalty Interpretation).
The EPO objective equals $V_{\lambda,\beta}^{\pi}=V^{\pi}+\lambda\mathcal{H}(\pi)-\rho_{k}\cdot\mathcal{P}(\pi)$, where $\mathcal{P}(\pi)$ is the exact penalty for violating the entropy corridor $\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$, and $\rho_{k}=\lambda\beta_{k}\alpha$ is an adaptive penalty weight.
This connects EPO to classical optimization: the smoothing loss is precisely the penalty for constraint violation, and the adaptive schedule implements curriculum-style enforcement—stronger early to guide entropy toward the corridor, relaxing later as stability is learned. While 6.1 establishes EPO’s principled foundation, it does not explain why EPO’s benefits are disproportionately large in multi-turn settings. The key insight is how entropy errors propagate:
Proposition 6.2 (Multi-Turn Entropy Stability).
Define cumulative entropy deviation $\Gamma_{T}(\pi):=\sum_{t=1}^{T}|\bar{\mathcal{H}}_{t}(\pi)-\bar{\mathcal{H}}^{ref}|$ over $T$ turns.
EPO: Under the corridor constraint, $\Gamma_{T}(\pi_{\theta})=O(T)$.
Standard entropy regularization: Without corridor enforcement, $\Gamma_{T}(\pi)=O(T^{2})$ in the worst case.
Without the corridor, entropy drift at turn $t$ affects turn $t+1$’s context, which affects turn t+2t+2—errors compound quadratically. EPO’s corridor acts as a guardrail, correcting deviations immediately. The advantage grows with dialogue length: $5.5\times$ for $T=10$, $10.5\times$ for $T=20$. The $O(T^{2})$ worst-case arises under conditions common in LLM training: sparse end-of-dialogue rewards, entropy collapse dynamics, and autoregressive context dependence (A.15).
Why Stability Improves Convergence. The $O(T)$ vs $O(T^{2})$ gap translates to convergence through two mechanisms: (i) gradient variance: entropy collapse causes extreme importance weights, increasing gradient noise; and (ii) trajectory diversity: collapsed entropy reduces the trajectories that reach informative rewards. We validate empirically that EPO achieves lower gradient variance and faster convergence (Appendix 5).
Proposition 6.3 (Performance Bound).
Under mild assumptions (Appendix A), EPO achieves:
$\small V^{\pi^{*}}-V^{\pi_{\theta}}\leq\underbrace{\frac{|\mathcal{D}|^{2}\epsilon^{2}}{2\lambda(1-\alpha\beta_{k})C}}_{\text{Optimization}}+\underbrace{\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}|^{1/H}}}_{\text{Entropy Bias}}+\underbrace{\lambda\beta_{k}\Delta^{\text{smooth}}_{\pi^{*}}}_{\text{Stability Cost}}$ (7)
The bound decomposes into optimization error (decreases with training), entropy bias (standard in entropy-regularized RL), and stability cost (price for corridor enforcement). Crucially, this cost often vanishes:
Proposition 6.4 (Zero Stability Cost).
If $\pi^{*}$ has moderate entropy ($\mathcal{H}(\pi^{*}|s)\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$), then $\Delta^{\text{smooth}}_{\pi^{*}}=0$—EPO matches standard bounds while providing stability guarantees for free.
In sparse-reward multi-turn tasks, optimal policies typically require moderate entropy to discover good trajectories, naturally satisfying this condition. Thus EPO provides stability "for free" precisely when cascade failure threatens most.
@@ 7 Conclusion
In this work, we identified and addressed the exploration-exploitation cascade failure, a fundamental challenge unique to training LLM agents in multi-turn environments with sparse rewards. Our proposed EPO framework addresses this through three synergistic mechanisms: trajectory-aware entropy computation, entropy smoothing regularization, and adaptive phase-based weighting. Together, these components prevent the severe entropy oscillations that destabilize training under shared policy parameters. Empirical results demonstrate up to 152% performance improvement on ScienceWorld and 19.8% on ALFWorld, transforming previously unstable sparse-reward scenarios into smoothly converging optimization. This work establishes that multi-turn LLM agent training requires fundamentally different entropy control than traditional RL, opening new directions for effective training methods for agentic LLMs.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning, specifically improving the training stability of LLM agents in multi-turn environments. Our method enables more reliable training of agents for beneficial applications such as scientific experimentation, household task assistance, and interactive problem-solving.
As with any advancement in autonomous agent capabilities, there exists potential for misuse if deployed without appropriate safeguards. However, our contribution is methodological in nature, focusing on training dynamics rather than enabling fundamentally new capabilities. We encourage practitioners to apply standard responsible AI practices when deploying agents trained with our method.
There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here beyond the considerations common to research on LLM agents.
Appendix A Theoretical Analysis: Full Proofs
A.1 Notation and Setup
We consider multi-turn dialogue generation as a finite-horizon MDP. Let $s_{0}$ denote the initial query, $H$ the maximum horizon (total tokens), and $T$ the number of dialogue turns. Each turn $t\in[T]$ generates $L_{t}$ tokens. The policy $\pi_{\theta}(a|s)$ maps states (conversation history) to distributions over tokens (actions).
Standard Entropy-Regularized Objective:
$V_{\lambda}^{\pi}(s_{0})=\mathbb{E}_{\pi}\left[\sum_{h=0}^{H-1}r_{h}+\lambda\mathcal{H}(\pi(\cdot|s_{h}))\;\Big|\;s_{0}\right]$ (8)
EPO Objective:
$V_{\lambda,\beta}^{\pi}(s_{0})=V_{\lambda}^{\pi}(s_{0})-\lambda\beta_{k}\cdot L^{\text{smooth}}_{\bar{\mathcal{H}}^{ref}}(\theta)$ (9)
where the smoothing loss is:
$L^{\text{smooth}}_{\bar{\mathcal{H}}^{ref}}(\theta)=\frac{\alpha}{|\mathcal{B}|}\sum_{(s,a)\in\mathcal{B}}\mathcal{P}(\mathcal{H}(\pi_{\theta}(\cdot|s));\bar{\mathcal{H}}^{ref})$ (10)
with corridor penalty:
$\mathcal{P}(\mathcal{H};\bar{\mathcal{H}}^{ref})=[\kappa_{l}\bar{\mathcal{H}}^{ref}-\mathcal{H}]_{+}+[\mathcal{H}-\kappa_{r}\bar{\mathcal{H}}^{ref}]_{+}$ (11)
A.2 Assumptions
Assumption A.1 (Fixed Entropy Reference).
The smoothing loss is evaluated against a fixed reference $\bar{\mathcal{H}}^{ref}>0$, representing the target entropy level.
Assumption A.2 (Entropy Corridor Constraint).
The learned policy $\pi_{\theta}$ maintains entropy within the corridor:
$\kappa_{l}\bar{\mathcal{H}}^{ref}\leq\mathcal{H}(\pi_{\theta}(\cdot|s))\leq\kappa_{r}\bar{\mathcal{H}}^{ref}$
for all reachable states $s\in\mathcal{S}(s_{0})$.
Assumption A.3 (Policy Regularity).
@@ A.2 Assumptions
The policy gradient satisfies $\|\nabla V^{\pi_{\theta}}_{\lambda,\beta}(\mathcal{D})\|\leq\epsilon$ and $\min_{a,s}\pi_{\theta}(a|s)\geq\delta_{\pi}>0$.
Assumption A.4 (Bounded Smoothing Strength).
The penalty weight satisfies $\alpha\beta_{\max}<1$, where $\beta_{\max}=1+e^{0}=2$. Equivalently, $\alpha<0.5$.
Remark A.5 (Justification).
A.2 captures EPO’s design intent—the smoothing penalty enforces corridor compliance. We verify this empirically in Appendix 5. A.4 ensures the EPO objective remains concave; our experiments use $\alpha=0.1<0.5$.
A.3 Key Definitions
Definition A.6 (Stability Cost).
For any policy $\tilde{\pi}$:
@@ A.3 Key Definitions
$\Delta^{\text{smooth}}_{\tilde{\pi}}(s_{0}):=\mathbb{E}_{\tilde{\pi}}\left[L^{\text{smooth}}_{\bar{\mathcal{H}}^{ref}}(\tilde{\theta})\;\Big|\;s_{0}\right]\geq 0$
Under A.2, $\Delta^{\text{smooth}}_{\pi_{\theta}}(s_{0})=0$.
Definition A.7 (Concentrability Coefficient).
For initial state $s_{0}$ and policy $\pi_{\theta}$:
$C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})=C^{2}_{d}\cdot\min_{h,s\in\mathcal{S}(s_{0})}\mathbb{P}^{\pi_{\theta}}_{h}(s)\left(\min_{a}\pi_{\theta}(a|s)\right)^{2}\cdot\min_{h,s\in\mathcal{S}(s_{0})}\frac{\mathbb{P}^{\pi_{\theta}}_{h}(s)}{\mathbb{P}^{\pi^{*}_{\lambda,\beta}}_{h}(s|s_{0})}$ (12)
where $C_{d}=(|\mathcal{A}|H)^{-1/2}$ and $\mathbb{P}^{\pi}_{h}(s)$ is the state visitation probability at step $h$.
Definition A.8 (EPO-Optimal Policy).
$\pi_{\lambda,\beta}^{*}:=\arg\max_{\pi\in\Pi}V_{\lambda,\beta}^{\pi}(\mathcal{D})$
Under A.4, the objective is concave, so this is well-defined.
Definition A.9 (Cumulative Entropy Deviation).
For $T$-turn dialogue with average entropy $\bar{\mathcal{H}}_{t}$ at turn $t$:
$\Gamma_{T}(\pi):=\sum_{t=1}^{T}|\bar{\mathcal{H}}_{t}(\pi)-\bar{\mathcal{H}}^{ref}|$
A.4 Preliminary Lemmas
Lemma A.10 (Penalty Function Properties).
The corridor penalty $\mathcal{P}(\mathcal{H};\bar{\mathcal{H}}^{ref})=[\kappa_{l}\bar{\mathcal{H}}^{ref}-\mathcal{H}]_{+}+[\mathcal{H}-\kappa_{r}\bar{\mathcal{H}}^{ref}]_{+}$ satisfies:
  1. 1.
    Non-negativity: $\mathcal{P}\geq 0$
  2. 2.
    Corridor property: $\mathcal{P}=0$ if and only if $\mathcal{H}\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$
  3. 3.
    Lipschitz continuity: $|\mathcal{P}(\mathcal{H}_{1})-\mathcal{P}(\mathcal{H}_{2})|\leq|\mathcal{H}_{1}-\mathcal{H}_{2}|$
@@ A.4 Preliminary Lemmas
Properties (1) and (2) follow from the definition of $\cdot]_{+}=\max(0,\cdot)$. For (3), both $\kappa_{l}\bar{\mathcal{H}}^{ref}-\mathcal{H}]_{+}$ and $\mathcal{H}-\kappa_{r}\bar{\mathcal{H}}^{ref}]_{+}$ are Lipschitz with constant 1, and at most one is active at any $\mathcal{H}$. ∎
Lemma A.11 (Gradient Towards Corridor).
For entropy outside the corridor:
$\nabla_{\theta}L^{\text{smooth}}=\begin{cases}-\alpha\nabla_{\theta}\mathcal{H}(\pi_{\theta})&\text{if }\mathcal{H}<\kappa_{l}\bar{\mathcal{H}}^{ref}\quad\text{(increases entropy)}\\ 0&\text{if }\mathcal{H}\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]\\ +\alpha\nabla_{\theta}\mathcal{H}(\pi_{\theta})&\text{if }\mathcal{H}>\kappa_{r}\bar{\mathcal{H}}^{ref}\quad\text{(decreases entropy)}\end{cases}$
Proof.
Taking gradients of $\mathcal{P}(\mathcal{H};\bar{\mathcal{H}}^{ref})$:
$\nabla_{\theta}\mathcal{P}=\begin{cases}-\nabla_{\theta}\mathcal{H}&\text{if }\mathcal{H}<\kappa_{l}\bar{\mathcal{H}}^{ref}\\ 0&\text{if }\mathcal{H}\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]\\ +\nabla_{\theta}\mathcal{H}&\text{if }\mathcal{H}>\kappa_{r}\bar{\mathcal{H}}^{ref}\end{cases}$
The result follows from $L^{\text{smooth}}=\alpha\cdot\mathbb{E}[\mathcal{P}]$. ∎
Lemma A.12 (Performance Difference (kakade2002approximately)).
For any policies $\pi,\pi^{\prime}$:
$V_{h}^{\pi}(s)-V_{h}^{\pi^{\prime}}(s)=\mathbb{E}_{\pi}\left[\sum_{t=h}^{H-1}A_{t}^{\pi^{\prime}}(s_{t},a_{t})\;\Big|\;s_{h}=s\right]$
where $A_{t}^{\pi^{\prime}}(s,a)=Q_{t}^{\pi^{\prime}}(s,a)-V_{t}^{\pi^{\prime}}(s)$ is the advantage function.
Lemma A.13 (Entropy Bias (shen2025entropy)).
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})+\lambda H\log\frac{|\mathcal{A}|}{|\mathcal{A}_{H}^{*}(s_{0})|^{1/H}}$
where $|\mathcal{A}_{H}^{*}(s_{0})|$ is the number of optimal action sequences from $s_{0}$.
A.5 Main Results
A.5.1 Exact Penalty Interpretation
See 6.1
Proof.
Expanding the EPO objective:
$V_{\lambda,\beta}^{\pi}$ $=V^{\pi}+\lambda\mathcal{H}(\pi)-\lambda\beta_{k}\cdot L^{\text{smooth}}_{\bar{\mathcal{H}}^{ref}}$ (13)
$=V^{\pi}+\lambda\mathcal{H}(\pi)-\lambda\beta_{k}\cdot\frac{\alpha}{|\mathcal{B}|}\sum_{(s,a)\in\mathcal{B}}\mathcal{P}(\mathcal{H}(\pi|s);\bar{\mathcal{H}}^{ref})$ (14)
$=V^{\pi}+\lambda\mathcal{H}(\pi)-\underbrace{\lambda\beta_{k}\alpha}_{\rho_{k}}\cdot\underbrace{\mathbb{E}_{s\sim\pi}[\mathcal{P}(\mathcal{H}(\pi|s);\bar{\mathcal{H}}^{ref})]}_{\mathcal{P}(\pi)}$ (15)
This matches the exact penalty formulation for the constraint $\mathcal{H}(\pi|s)\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$. ∎
Remark A.14 (Connection to Classical Optimization).
The exact penalty method (kakade2002approximately) guarantees that for sufficiently large $\rho$, the unconstrained optimum coincides with the constrained optimum. EPO’s adaptive schedule $\rho_{k}=\lambda\beta_{k}\alpha$ provides curriculum-style enforcement: weaker early (exploration), stronger later (exploitation).
A.5.2 Multi-Turn Stability
See 6.2
Proof.
Part (i): EPO bound.
Under A.2, for all turns $t$ and states $s$:
$\mathcal{H}(\pi_{\theta}(\cdot|s))\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$
The average entropy at turn $t$ satisfies:
$\bar{\mathcal{H}}_{t}(\pi_{\theta})=\frac{1}{L_{t}}\sum_{h\in\text{turn }t}\mathcal{H}(\pi_{\theta}(\cdot|s_{h}))\in[\kappa_{l}\bar{\mathcal{H}}^{ref},\kappa_{r}\bar{\mathcal{H}}^{ref}]$
Therefore:
$|\bar{\mathcal{H}}_{t}(\pi_{\theta})-\bar{\mathcal{H}}^{ref}|\leq\max\{|1-\kappa_{l}|,|\kappa_{r}-1|\}\cdot\bar{\mathcal{H}}^{ref}\leq\frac{\kappa_{r}-\kappa_{l}}{2}\cdot\bar{\mathcal{H}}^{ref}:=\Delta_{\max}$
Summing over $T$ turns:
$\Gamma_{T}(\pi_{\theta})=\sum_{t=1}^{T}|\bar{\mathcal{H}}_{t}(\pi_{\theta})-\bar{\mathcal{H}}^{ref}|\leq T\cdot\Delta_{\max}=O(T)$
Part (ii): Entropy regularization worst case.
Without corridor enforcement, entropy at turn $t$ depends on context from turns $1,\ldots,t-1$. Consider the cascade failure scenario where each turn introduces drift $\delta$:
$\bar{\mathcal{H}}_{t}-\bar{\mathcal{H}}^{ref}=\sum_{s=1}^{t}\delta_{s}$
If drifts are systematic (e.g., consistently decreasing entropy), $|\bar{\mathcal{H}}_{t}-\bar{\mathcal{H}}^{ref}|\leq t\cdot\delta_{\max}$.
Summing:
$\Gamma_{T}(\pi)\leq\sum_{t=1}^{T}t\cdot\delta_{\max}=\frac{T(T+1)}{2}\cdot\delta_{\max}=O(T^{2})$
Remark A.15 (Conditions for Cascade Failure).
The $O(T^{2})$ worst-case arises under specific but common conditions in multi-turn LLM training:
(i) Sparse rewards: When rewards are only observed at dialogue end (e.g., task success/failure), there is no per-turn gradient signal to correct entropy drift early. Each turn’s drift goes unchecked until final reward, by which point deviations have compounded.
(ii) Entropy collapse dynamics: Once entropy begins decreasing, the softmax policy assigns increasingly extreme probabilities. This reduces exploration, which reduces trajectory diversity, which reduces the probability of discovering reward signal, which further reduces entropy—a positive feedback loop.
(iii) Context-dependent amplification: In autoregressive generation, turn $t$’s output becomes turn $t+1$’s context. Low-entropy (confident) outputs at turn $t$ bias the model toward similar low-entropy behavior at turn $t+1$, causing drift to accumulate rather than cancel.
These conditions are prevalent in multi-turn LLM training, so the $O(T^{2})$ bound captures realistic failure modes rather than pathological edge cases.
Remark A.16 (Practical Implication).
The $O(T)$ vs $O(T^{2})$ gap implies EPO’s advantage grows with dialogue length. For $T=10$ turns, the worst-case ratio is $\frac{T(T+1)/2}{T}=\frac{T+1}{2}=5.5\times$. For $T=20$, this becomes $10.5\times$. This explains why EPO’s empirical benefits are most pronounced in long multi-turn tasks (Appendix 5).
A.5.3 Connection Between Entropy Stability and Convergence
Remark A.17 (Why Entropy Stability Improves Convergence).
While we do not formally prove a bound relating $\Gamma_{T}$ to suboptimality, we identify two mechanisms through which entropy stability improves convergence:
Mechanism 1: Importance Weight Variance. Policy gradient estimators use importance weights $w=\pi_{\theta}(a|s)/\pi_{\text{old}}(a|s)$. When entropy collapses ($\mathcal{H}\to 0$), policies become near-deterministic, causing $w\to\infty$ for some actions and $w\to 0$ for others. This increases gradient variance, slowing convergence. EPO’s corridor constraint bounds entropy away from zero, limiting weight variance.
Mechanism 2: Trajectory Diversity. In multi-turn generation with sparse rewards, gradient signal comes from trajectories that reach reward. Collapsed entropy reduces trajectory diversity exponentially with horizon: if per-step entropy drops by factor $c<1$, trajectory diversity drops by $c^{T}$. The corridor constraint maintains per-step entropy, preserving trajectory diversity throughout training.
Empirical Validation. Rather than formalizing these mechanisms into a bound with unverifiable constants, we directly measure the effect in Appendix 5: EPO achieves lower gradient variance and faster convergence compared to standard entropy regularization, consistent with this intuition.
A.5.4 Performance Bound
See 6.3
Proof.
Step 1: Decomposition through EPO-optimal policy.
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})=\underbrace{[V^{\pi^{*}}-V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}]}_{(I)}+\underbrace{[V_{\lambda,\beta}^{\pi_{\lambda,\beta}^{*}}-V_{\lambda,\beta}^{\pi_{\theta}}]}_{(II)}+\underbrace{[V_{\lambda,\beta}^{\pi_{\theta}}-V^{\pi_{\theta}}]}_{(III)}$ (16)
Step 2: Bounding Term (II).
Under A.4, the EPO objective has effective entropy coefficient:
$\lambda_{\text{eff}}=\lambda(1-\alpha\beta_{k})>0$
Since $\alpha\beta_{k}<1$, the objective $V_{\lambda,\beta}$ remains concave in policy parameters. The gradient-to-performance bound from shen2025entropy applies:
$(II)\leq\frac{|\mathcal{D}|^{2}}{2\lambda_{\text{eff}}\cdot C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})}\epsilon^{2}=\frac{|\mathcal{D}|^{2}\epsilon^{2}}{2\lambda(1-\alpha\beta_{k})C_{\lambda,\beta}^{\pi_{\theta}}(s_{0})}$ (17)
Step 3: Bounding Term (III).
Under A.2, $L^{\text{smooth}}(\theta)=0$ by A.10(2). Thus:
$(III)=V_{\lambda,\beta}^{\pi_{\theta}}(s_{0})-V^{\pi_{\theta}}(s_{0})=\lambda\mathcal{H}(\pi_{\theta}|s_{0})-0=\lambda\mathcal{H}(\pi_{\theta}|s_{0})$ (18)
Step 4: Bounding Term (I).
By definition of $\pi_{\lambda,\beta}^{*}$ as the maximizer: